# Doubly-charmed pentaquark states in a mass splitting model

### v1

@@ I Introduction
The intriguing experimental findings also triggered discussions on possible doubly-charmed pentaquark states Zhou:2018bkn ; Wang:2018lhz ; Park:2018oib ; Giannuzzi:2019esi ; Xing:2021yid ; Andreev:2022qdu ; Ozdem:2022vip ; Park:2023ygm ; Guo:2017vcf ; Chen:2017vai ; Dias:2018qhp ; Shimizu:2017xrg ; Yang:2020twg ; Chen:2021kad ; Zhou:2022gra ; Wang:2022aga ; Shen:2022zvd ; Yalikun:2023waw ; Wang:2023mdj ; Wang:2023eng ; Liu:2023clr ; Duan:2024uuf ; Wang:2024brl . Replacing the light quark in the conventional baryon $\Xi_{cc}$ with two light antiquarks, one gets a doubly-charmed tetraquark $T_{cc}$ which was predicted in various theoretical models long time ago Ballot:1983iv ; Zouzou:1986qh . Replacing one light antiquark in the $T_{cc}$ with two quarks further, or putting a $q\bar{q}$ pair inside the $\Xi_{cc}$, one gets a potential doubly-charmed pentaquark state. To date, no such experimental candidate has been announced. However, the existence of such states cannot be excluded. Insights may be drawn from the search for $\Xi_{cc}$ and $T_{cc}$. Two decades ago, the SELEX Collaboration claimed the observation of the quark-model-predicted doubly-charmed baryon $\Xi_{cc}^{+}$ in the $\Lambda_{c}^{+}K^{-}\pi^{+}$ three-body decay channel SELEX:2002wqn , but other collaborations did not confirm its existence. In 2017, its partner state $\Xi_{cc}^{++}$ was observed by the LHCb LHCb:2017iph in the $\Lambda_{c}^{+}K^{-}\pi^{+}\pi^{+}$ channel with a $\sim$100 MeV higher mass than the SELEX $\Xi_{cc}^{+}$. The LHCb also observed $\Xi_{cc}^{++}$ in the channel $\Xi_{c}^{+}\pi^{+}$ LHCb:2018pcs . Later in 2022, the LHCb succeeded in finding a doubly-charmed four-quark state, $T_{cc}^{+}(3875)$, in the $D^{0}D^{0}\pi^{+}$ mass distribution just below the $D^{*+}D^{0}$ threshold LHCb:2021auc ; LHCb:2021vvq . The observations of $\Xi_{cc}^{++}$ and $T_{cc}^{+}(3875)$ lend credibility to find the doubly-charmed pentaquarks finally, if they should exist from a theoretical point of view.
@@
In the literature, a number of discussions on $QQqq\bar{q}$ ($Q=c,b$; $q=u,d,s$) with respect to their spectra, decays, electromagnetic properties, and so on in configurations such as compact pentaquark Zhou:2018bkn ; Wang:2018lhz ; Park:2018oib ; Giannuzzi:2019esi ; Xing:2021yid ; Andreev:2022qdu ; Ozdem:2022vip ; Park:2023ygm and hadronic molecule Guo:2017vcf ; Chen:2017vai ; Dias:2018qhp ; Shimizu:2017xrg ; Yang:2020twg ; Chen:2021kad ; Zhou:2022gra ; Wang:2022aga ; Shen:2022zvd ; Yalikun:2023waw ; Wang:2023mdj ; Wang:2023eng ; Liu:2023clr ; Duan:2024uuf ; Wang:2024brl ; Xu:2010fc ; Sheng:2024hkf can be found. The electromagnetic properties were studied in Refs. Zhou:2022gra ; Wang:2023ael ; Ozdem:2024yel ; Ozdem:2022vip . In Refs. Zhou:2018bkn ; Park:2018oib ; Park:2023ygm , the authors explored spectra in the compact pentaquark picture utilizing color-magnetic interaction (CMI) models, while QCD sum rules were adopted for the study of compact states in Ref. Wang:2018lhz . A quark potential model was employed in a diquark-diquark-antiquark configration to estimate the masses in Ref. Giannuzzi:2019esi . The authors of Ref. Xing:2021yid obtained spectra of $ccqq\bar{q}$ states in a triquark-diquark configuration and calculated decay widths using the operator product expansion technique. Compared with the compact configuration, the baryon-meson molecule is a more popular one in exploring doubly-heavy pentaquark-like states. Such studies were performed with various methods, like one-boson-exchange potential model Shimizu:2017xrg ; Chen:2017vai ; Dias:2018qhp ; Chen:2021kad ; Dong:2021bvy ; Shen:2022zvd ; Wang:2023aob ; Yalikun:2023waw ; Xu:2010fc , chiral quark model Yang:2020twg , chiral effective field theory Xu:2010fc ; Guo:2017vcf ; Chen:2021htr , QCD sum rules Duan:2024uuf ; Wang:2024brl , local hidden gauge approach Wang:2022aga ; Wang:2023mdj , resonating group method based on the QDCSM framework Liu:2023clr , near-threshold interaction with the light $qq$ and $q\bar{q}$ correlations Wang:2023eng , etc.
@@ II.1 Mass formulae
Here, the subscript $i$ represents the $i$th constituent quark. The effective quark mass, denoted as $m_{i}$, contains contributions from various terms such as quark kinetic energy, color Coulomb interaction, and color confinement. The parameter $C_{ij}$ is the effective coupling constant characterizing the interaction strength between the $i$th and $j$th quark components. One extracts these parameters from the measured masses of conventional hadrons. $\lambda_{i}$ and $\sigma_{i}$ denote the $SU(3)$ Gell-Mann matrices and $SU(2)$ Pauli matrices for the $i$th quark, respectively. For the antiquark, $\lambda_{i}$ should be replaced by $-\lambda_{i}^{*}$. After one establishes the spin-color base vectors, the CMI matrix $\langle H_{CMI}\rangle$ and then its eigenvalues $E_{CMI}$’s and eigenvectors can be determined. It is easy to write down the mass formula corresponding to Eq. (1)
@@ II.2 Spin-color base vectors
The calculations of $\langle H_{CMI}\rangle$, $E_{CMI}$, and rearrangement decay widths need the flavor-spin-color wave functions that are constrained by the Pauli principle. We do not consider the isospin breaking effects and the systems to be studied are $ccnn\bar{n}$, $ccnn\bar{s}$, $ccss\bar{n}$, $ccss\bar{s}$, $ccns\bar{n}$, and $ccns\bar{s}$.
@@
$\chi_{SS}$ ::: $\left\{\begin{array}[]{ccc}\chi_{1}&=&[((cc)^{1}(qq)^{1})^{2}\bar{q}]^{\frac{5}{2}},\\ \chi_{2}&=&[((cc)^{1}(qq)^{1})^{2}\bar{q}]^{\frac{3}{2}},\\ \chi_{3}&=&[((cc)^{1}(qq)^{1})^{1}\bar{q}]^{\frac{3}{2}},\\ \chi_{4}&=&[((cc)^{1}(qq)^{1})^{0}\bar{q}]^{\frac{1}{2}},\\ \chi_{5}&=&[((cc)^{1}(qq)^{1})^{1}\bar{q}]^{\frac{1}{2}},\end{array}\right.$ (12)
$\chi_{SA}$ ::: $\left\{\begin{array}[]{ccc}\chi_{6}&=&[((cc)^{1}(qq)^{0})^{1}\bar{q}]^{\frac{3}{2}},\\ \chi_{7}&=&[((cc)^{1}(qq)^{0})^{1}\bar{q}]^{\frac{1}{2}},\end{array}\right.$ (15)
$\chi_{AS}$ ::: $\left\{\begin{array}[]{ccc}\chi_{8}&=&[((cc)^{0}(qq)^{1})^{1}\bar{q}]^{\frac{3}{2}},\\ \chi_{9}&=&[((cc)^{0}(qq)^{1})^{1}\bar{q}]^{\frac{1}{2}},\end{array}\right.$ (18)
$\chi_{AA}$ ::: $\chi_{10}=[((cc)^{0}(qq)^{0})^{0}\bar{q}]^{\frac{1}{2}},$
@@ III.1 Model parameters
Utilizing the aforementioned parameters, we can derive the mass spectra and rearrangement decays of $ccqq\bar{q}$ compact pentaquark states. We classify the studied systems into eight cases, $(ccnn\bar{n})^{I_{nn}=1}$, $(ccnn\bar{n})^{I_{nn}=0}$, $(ccnn\bar{s})^{I_{nn}=1}$, $(ccnn\bar{s})^{I_{nn}=0}$, $ccss\bar{n}$, $ccss\bar{s}$, $ccns\bar{n}$, and $ccns\bar{s}$. It should be noted that the $(ccnn\bar{n})^{I_{nn}=1}$ case has degenerate $I=\frac{3}{2}$ and $\frac{1}{2}$ states while the $ccns\bar{n}$ case has degenerate $I=1$ and 0 states since the isospin breaking effects are not considered. For decays of degenerate states with different isospins but same spin, the partial widths of the same channel may be different due to their flavor wavefunctions, which can be seen in the following discussions.
@@ III.2 Spectra and decay widths of $ccnn\bar{n}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}165.6\end{array}\right)$ $\left(\begin{array}[]{c}165.6\end{array}\right)$ [$\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4548.2\end{array}\right)$ $\left(\begin{array}[]{c}4585.6\end{array}\right)$ $\left(\begin{array}[]{c}4437.0\end{array}\right)$ $\left(\begin{array}[]{c}4699.2\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-77.7&69.5&267.4&59.0\\ 69.5&95.6&-87.6&-5.6\\ 267.4&-87.6&73.3&-16.0\\ 59.0&-5.6&-16.0&22.0\end{array}\right)$ $\left(\begin{array}[]{c}282.3\\ 121.7\\ 27.6\\ -318.4\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.57,0.16,-0.80,-0.08\}\\ \{-0.31,-0.94,0.04,-0.14\}\\ \{0.03,-0.16,-0.16,0.97\}\\ \{0.76,-0.25,-0.58,-0.16\}\end{array}\right]$ $\left(\begin{array}[]{c}4664.9\\ 4504.3\\ 4410.2\\ 4064.2\end{array}\right)$ $\left(\begin{array}[]{c}4702.3\\ 4541.7\\ 4447.6\\ 4101.6\end{array}\right)$ $\left(\begin{array}[]{c}4553.7\\ 4393.1\\ 4299.0\\ 3953.1\end{array}\right)$ $\left(\begin{array}[]{c}4815.9\\ 4655.3\\ 4561.2\\ 4215.2\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}36.3&87.9&-169.1&-37.3\\ 87.9&-50.4&271.2&-84.8\\ -169.1&271.2&99.7&-16.0\\ -37.3&-84.8&-16.0&141.6\end{array}\right)$ $\left(\begin{array}[]{c}343.7\\ 189.7\\ 34.4\\ -340.6\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.25,0.53,0.78,-0.24\}\\ \{-0.56,-0.23,0.22,0.77\}\\ \{-0.67,-0.43,-0.10,-0.59\}\\ \{0.42,-0.69,0.58,-0.07\}\end{array}\right]$ $\left(\begin{array}[]{c}4726.3\\ 4572.3\\ 4417.0\\ 4042.0\end{array}\right)$ $\left(\begin{array}[]{c}4763.7\\ 4609.7\\ 4454.4\\ 4079.4\end{array}\right)$ $\left(\begin{array}[]{c}4615.1\\ 4461.1\\ 4305.8\\ 3930.9\end{array}\right)$ $\left(\begin{array}[]{c}4877.3\\ 4723.3\\ 4568.0\\ 4193.0\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}201.2\end{array}\right)$ $\left(\begin{array}[]{c}201.2\end{array}\right)$ [$\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4583.8\end{array}\right)$ $\left(\begin{array}[]{c}4621.2\end{array}\right)$ $\left(\begin{array}[]{c}4472.6\end{array}\right)$ $\left(\begin{array}[]{c}4734.8\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccc}-275.1&232.7&267.4\\ 232.7&52.6&-87.6\\ 267.4&-87.6&-121.1\end{array}\right)$ $\left(\begin{array}[]{c}185.2\\ 40.3\\ -569.1\end{array}\right)$ [$\left[\begin{array}[]{ccc}\{0.55,0.80,0.25\}\\ \{0.33,-0.48,0.81\}\\ \{0.77,-0.36,-0.53\}\end{array}\right]$ $\left(\begin{array}[]{c}4567.8\\ 4422.9\\ 3813.5\end{array}\right)$ $\left(\begin{array}[]{c}4605.2\\ 4460.3\\ 3850.9\end{array}\right)$ $\left(\begin{array}[]{c}4456.7\\ 4311.7\\ 3702.3\end{array}\right)$ $\left(\begin{array}[]{c}4718.8\\ 4573.9\\ 3964.5\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}-69.3&27.7&294.3&-169.1\\ 27.7&-134.4&0.0&64.7\\ 294.3&0.0&-233.2&271.2\\ -169.1&64.7&271.2&-173.9\end{array}\right)$ $\left(\begin{array}[]{c}162.8\\ 37.2\\ -142.3\\ -668.4\end{array}\right)$ [$\left[\begin{array}[]{cccc}\{-0.71,-0.11,-0.67,-0.20\}\\ \{-0.50,0.22,0.26,0.80\}\\ \{-0.08,-0.96,0.20,0.15\}\\ \{0.49,-0.09,-0.67,0.55\}\end{array}\right]$ $\left(\begin{array}[]{c}4545.4\\ 4419.8\\ 4240.3\\ 3714.2\end{array}\right)$ $\left(\begin{array}[]{c}4582.8\\ 4457.2\\ 4277.7\\ 3751.6\end{array}\right)$ $\left(\begin{array}[]{c}4434.3\\ 4308.6\\ 4129.1\\ 3603.0\end{array}\right)$ $\left(\begin{array}[]{c}4696.4\\ 4570.8\\ 4391.3\\ 3865.2\end{array}\right)$

### v2

@@ I Introduction
The intriguing experimental findings also triggered discussions on possible doubly-charmed pentaquark states Zhou:2018bkn ; Wang:2018lhz ; Park:2018oib ; Giannuzzi:2019esi ; Xing:2021yid ; Andreev:2022qdu ; Ozdem:2022vip ; Park:2023ygm ; Yang:2024okq ; Guo:2017vcf ; Chen:2017vai ; Dias:2018qhp ; Yan:2018zdt ; Shimizu:2017xrg ; Yang:2020twg ; Chen:2021kad ; Zhou:2022gra ; Wang:2022aga ; Shen:2022zvd ; Yalikun:2023waw ; Wang:2023mdj ; Wang:2023eng ; Liu:2023clr ; Duan:2024uuf ; Wang:2024brl . Replacing the light quark in the conventional baryon $\Xi_{cc}$ with two light antiquarks, one gets a doubly-charmed tetraquark $T_{cc}$ which was predicted in various theoretical models long time ago Ballot:1983iv ; Zouzou:1986qh . Replacing one light antiquark in the $T_{cc}$ with two quarks further, or putting a $q\bar{q}$ pair inside the $\Xi_{cc}$, one gets a potential doubly-charmed pentaquark state. To date, no such experimental candidate has been announced. However, the existence of such states cannot be excluded. Insights may be drawn from the search for $\Xi_{cc}$ and $T_{cc}$. Two decades ago, the SELEX Collaboration claimed the observation of the quark-model-predicted doubly-charmed baryon $\Xi_{cc}^{+}$ in the $\Lambda_{c}^{+}K^{-}\pi^{+}$ three-body decay channel SELEX:2002wqn , but other collaborations did not confirm its existence. In 2017, its partner state $\Xi_{cc}^{++}$ was observed by the LHCb LHCb:2017iph in the $\Lambda_{c}^{+}K^{-}\pi^{+}\pi^{+}$ channel with a $\sim$100 MeV higher mass than the SELEX $\Xi_{cc}^{+}$. The LHCb also observed $\Xi_{cc}^{++}$ in the channel $\Xi_{c}^{+}\pi^{+}$ LHCb:2018pcs . Later in 2022, the LHCb succeeded in finding a doubly-charmed four-quark state, $T_{cc}^{+}(3875)$, in the $D^{0}D^{0}\pi^{+}$ mass distribution just below the $D^{*+}D^{0}$ threshold LHCb:2021auc ; LHCb:2021vvq . The observations of $\Xi_{cc}^{++}$ and $T_{cc}^{+}(3875)$ lend credibility to find the doubly-charmed pentaquarks finally, if they should exist from a theoretical point of view.
@@
In the literature, a number of discussions on $QQqq\bar{q}$ ($Q=c,b$; $q=u,d,s$) with respect to their spectra, decays, electromagnetic properties, and so on in configurations such as compact pentaquark Zhou:2018bkn ; Wang:2018lhz ; Park:2018oib ; Giannuzzi:2019esi ; Xing:2021yid ; Andreev:2022qdu ; Ozdem:2022vip ; Park:2023ygm ; Yang:2024okq and hadronic molecule Guo:2017vcf ; Chen:2017vai ; Dias:2018qhp ; Yan:2018zdt ; Shimizu:2017xrg ; Yang:2020twg ; Chen:2021kad ; Zhou:2022gra ; Wang:2022aga ; Shen:2022zvd ; Yalikun:2023waw ; Wang:2023mdj ; Wang:2023eng ; Liu:2023clr ; Duan:2024uuf ; Wang:2024brl ; Xu:2010fc ; Sheng:2024hkf ; Yang:2024okq can be found. The electromagnetic properties were studied in Refs. Zhou:2022gra ; Wang:2023ael ; Ozdem:2024yel ; Ozdem:2022vip . In Refs. Zhou:2018bkn ; Park:2018oib ; Park:2023ygm , the authors explored spectra in the compact pentaquark picture utilizing color-magnetic interaction (CMI) models, while QCD sum rules were adopted for the study of compact states in Ref. Wang:2018lhz ; Yang:2024okq . A quark potential model was employed in a diquark-diquark-antiquark configration to estimate the masses in Ref. Giannuzzi:2019esi . The authors of Ref. Xing:2021yid obtained spectra of $ccqq\bar{q}$ states in a triquark-diquark configuration and calculated decay widths using the operator product expansion technique. Compared with the compact configuration, the baryon-meson molecule is a more popular one in exploring doubly-heavy pentaquark-like states. Such studies were performed with various methods, like one-boson-exchange potential model Shimizu:2017xrg ; Chen:2017vai ; Dias:2018qhp ; Chen:2021kad ; Dong:2021bvy ; Shen:2022zvd ; Wang:2023aob ; Yalikun:2023waw ; Xu:2010fc , chiral quark model Yang:2020twg , chiral effective field theory Xu:2010fc ; Guo:2017vcf ; Yan:2018zdt ; Chen:2021htr , QCD sum rules Duan:2024uuf ; Wang:2024brl ; Yang:2024okq , local hidden gauge approach Wang:2022aga ; Wang:2023mdj , resonating group method based on the QDCSM framework Liu:2023clr , near-threshold interaction with the light $qq$ and $q\bar{q}$ correlations Wang:2023eng , etc.
@@ II.1 Mass formulae
Here, the subscript $i$ represents the $i$th constituent quark. The effective quark mass, denoted as $m_{i}$, contains contributions from various terms such as quark kinetic energy, color Coulomb interaction, and color confinement. The parameter $C_{ij}$ is the effective coupling constant characterizing the interaction strength between the $i$th and $j$th quark components. One extracts these parameters from the measured masses of conventional hadrons. $\lambda_{i}$ and $\sigma_{i}$ denote the $SU(3)$ Gell-Mann matrices and $SU(2)$ Pauli matrices for the $i$th quark, respectively. For the antiquark, $\lambda_{i}$ should be replaced by $\lambda_{i}^{*}$. After one establishes the spin-color base vectors, the CMI matrix $\langle H_{CMI}\rangle$ and then its eigenvalues $E_{CMI}$’s and eigenvectors can be determined. It is easy to write down the mass formula corresponding to Eq. (1)
@@ II.2 Spin-color base vectors
The calculations of $\langle H_{CMI}\rangle$, $E_{CMI}$, and rearrangement decay widths need the flavor-spin-color wave functions that are constrained by the Pauli principle. We do not consider the isospin breaking effects and the systems to be studied are $cnn\bar{n}, cc$$nn\bar{s}, cc$$s\bar{n}, ccss$$s\bar{s}, cc$$cns\bar{n}$, and $ccns\bar{s}$.
@@
$\chi_{SS}$ $:$ $\left\{\begin{array}[]{ccc}\chi_{1}&=&[((cc)^{1}(qq)^{1})^{2}\bar{q}]^{\frac{5}{2}},\\ \chi_{2}&=&[((cc)^{1}(qq)^{1})^{2}\bar{q}]^{\frac{3}{2}},\\ \chi_{3}&=&[((cc)^{1}(qq)^{1})^{1}\bar{q}]^{\frac{3}{2}},\\ \chi_{4}&=&[((cc)^{1}(qq)^{1})^{0}\bar{q}]^{\frac{1}{2}},\\ \chi_{5}&=&[((cc)^{1}(qq)^{1})^{1}\bar{q}]^{\frac{1}{2}},\end{array}\right$. (12)
$\chi_{SA}$ $:$ $\left\{\begin{array}[]{ccc}\chi_{6}&=&[((cc)^{1}(qq)^{0})^{1}\bar{q}]^{\frac{3}{2}},\\ \chi_{7}&=&[((cc)^{1}(qq)^{0})^{1}\bar{q}]^{\frac{1}{2}},\end{array}\right$. (15)
$\chi_{AS}$ $:$ $\left\{\begin{array}[]{ccc}\chi_{8}&=&[((cc)^{0}(qq)^{1})^{1}\bar{q}]^{\frac{3}{2}},\\ \chi_{9}&=&[((cc)^{0}(qq)^{1})^{1}\bar{q}]^{\frac{1}{2}},\end{array}\right$. (18)
$\chi_{AA}$ $:$ $\chi_{10}=[((cc)^{0}(qq)^{0})^{0}\bar{q}]^{\frac{1}{2}},$
@@ III.1 Model parameters
Utilizing the aforementioned parameters, we can derive the mass spectra and rearrangement decays of $ccqq\bar{q}$ compact pentaquark states. We classify the studied systems into eight cases, $(ccnn\bar{n})^{I_{nn}=1}$, $(ccnn\bar{n})^{I_{nn}=0}$, $(ccnn\bar{s})^{I_{nn}=1}$, $(ccnn\bar{s})^{I_{nn}=0}$, $\bar{n}, ccsss¯ccs$$\bar{s}$, $ccns\bar{n}$, and $ccns\bar{s}$. It should be noted that the $(ccnn\bar{n})^{I_{nn}=1}$ case has degenerate $I=\frac{3}{2}$ and $\frac{1}{2}$ states while the $ccns\bar{n}$ case has degenerate $I=1$ and 0 states since the isospin breaking effects are not considered. For decays of degenerate states with different isospins but same spin, the partial widths of the same channel may be different due to their flavor wavefunctions, which can be seen in the following discussions.
@@ III.2 Spectra and decay widths of $ccnn\bar{n}$ states
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}165.6\end{array}\right)$ $\left(\begin{array}[]{c}165.6\end{array}\right)$ $\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4548.2\end{array}\right)$ $\left(\begin{array}[]{c}4585.6\end{array}\right)$ $\left(\begin{array}[]{c}4437.0\end{array}\right)$ $\left(\begin{array}[]{c}4699.2\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{cccc}-77.7&69.5&267.4&59.0\\ 69.5&95.6&-87.6&-5.6\\ 267.4&-87.6&73.3&-16.0\\ 59.0&-5.6&-16.0&22.0\end{array}\right)$ $\left(\begin{array}[]{c}282.3\\ 121.7\\ 27.6\\ -318.4\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.57,0.16,-0.80,-0.08\}\\ \{-0.31,-0.94,0.04,-0.14\}\\ \{0.03,-0.16,-0.16,0.97\}\\ \{0.76,-0.25,-0.58,-0.16\}\end{array}\right]$ $\left(\begin{array}[]{c}4664.9\\ 4504.3\\ 4410.2\\ 4064.2\end{array}\right)$ $\left(\begin{array}[]{c}4702.3\\ 4541.7\\ 4447.6\\ 4101.6\end{array}\right)$ $\left(\begin{array}[]{c}4553.7\\ 4393.1\\ 4299.0\\ 3953.1\end{array}\right)$ $\left(\begin{array}[]{c}4815.9\\ 4655.3\\ 4561.2\\ 4215.2\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}36.3&87.9&-169.1&-37.3\\ 87.9&-50.4&271.2&-84.8\\ -169.1&271.2&99.7&-16.0\\ -37.3&-84.8&-16.0&141.6\end{array}\right)$ $\left(\begin{array}[]{c}343.7\\ 189.7\\ 34.4\\ -340.6\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.25,0.53,0.78,-0.24\}\\ \{-0.56,-0.23,0.22,0.77\}\\ \{-0.67,-0.43,-0.10,-0.59\}\\ \{0.42,-0.69,0.58,-0.07\}\end{array}\right]$ $\left(\begin{array}[]{c}4726.3\\ 4572.3\\ 4417.0\\ 4042.0\end{array}\right)$ $\left(\begin{array}[]{c}4763.7\\ 4609.7\\ 4454.4\\ 4079.4\end{array}\right)$ $\left(\begin{array}[]{c}4615.1\\ 4461.1\\ 4305.8\\ 3930.9\end{array}\right)$ $\left(\begin{array}[]{c}4877.3\\ 4723.3\\ 4568.0\\ 4193.0\end{array}\right)$
@@
$\frac{5}{2}^{-}$ $\left(\begin{array}[]{c}201.2\end{array}\right)$ $\left(\begin{array}[]{c}201.2\end{array}\right)$ $\left[\begin{array}[]{c}\{1.00\}\end{array}\right]$ $\left(\begin{array}[]{c}4583.8\end{array}\right)$ $\left(\begin{array}[]{c}4621.2\end{array}\right)$ $\left(\begin{array}[]{c}4472.6\end{array}\right)$ $\left(\begin{array}[]{c}4734.8\end{array}\right)$
$\frac{3}{2}^{-}$ $\left(\begin{array}[]{ccc}-275.1&232.7&267.4\\ 232.7&52.6&-87.6\\ 267.4&-87.6&-121.1\end{array}\right)$ $\left(\begin{array}[]{c}185.2\\ 40.3\\ -569.1\end{array}\right)$ $\left[\begin{array}[]{ccc}\{0.55,0.80,0.25\}\\ \{0.33,-0.48,0.81\}\\ \{0.77,-0.36,-0.53\}\end{array}\right]$ $\left(\begin{array}[]{c}4567.8\\ 4422.9\\ 3813.5\end{array}\right)$ $\left(\begin{array}[]{c}4605.2\\ 4460.3\\ 3850.9\end{array}\right)$ $\left(\begin{array}[]{c}4456.7\\ 4311.7\\ 3702.3\end{array}\right)$ $\left(\begin{array}[]{c}4718.8\\ 4573.9\\ 3964.5\end{array}\right)$
$\frac{1}{2}^{-}$ $\left(\begin{array}[]{cccc}-69.3&27.7&294.3&-169.1\\ 27.7&-134.4&0.0&64.7\\ 294.3&0.0&-233.2&271.2\\ -169.1&64.7&271.2&-173.9\end{array}\right)$ $\left(\begin{array}[]{c}162.8\\ 37.2\\ -142.3\\ -668.4\end{array}\right)$ $\left[\begin{array}[]{cccc}\{-0.71,-0.11,-0.67,-0.20\}\\ \{-0.50,0.22,0.26,0.80\}\\ \{-0.08,-0.96,0.20,0.15\}\\ \{0.49,-0.09,-0.67,0.55\}\end{array}\right]$ $\left(\begin{array}[]{c}4545.4\\ 4419.8\\ 4240.3\\ 3714.2\end{array}\right)$ $\left(\begin{array}[]{c}4582.8\\ 4457.2\\ 4277.7\\ 3751.6\end{array}\right)$ $\left(\begin{array}[]{c}4434.3\\ 4308.6\\ 4129.1\\ 3603.0\end{array}\right)$ $\left(\begin{array}[]{c}4696.4\\ 4570.8\\ 4391.3\\ 3865.2\end{array}\right)$